# Imperial College London, London, SW7 2AZ, UK

### v1

@@ Acknowledgements
We acknowledge interesting conversations with Ahmed Almheiri, Cliff Burgess, Thomas Colas, Paolo Creminelli, Steve Gratton, Gonzalo Villa, Greg Kaplanek, Simon Lin, Juan Maldacena, Daniel Mata-Pacheco, Mehrdad Mirbabayi, Shoy Ouseph, Pellegrino Piantadosi, Mario Ramos-Hamud and Luis Zapata. We are particularly thankful to Enrico Pajer for detailed comments on a previous version of this paper. SC thanks the hospitality of New York University Abu Dhabi where part of this work was carried. SC is supported in part by the STFC Consolidated Grants ST/T000791/1 and ST/X000575/1 and by a Simons Investigator award 690508. The research of FQ is funded by a NYUAD research grant.
@@ Appendix A Details on the path integral derivations
$S[\phi]=\int d^{4}x\sqrt{-g}\left(-(\partial_{\mu}\phi)^{2}+\mathcal{L}_{\mathrm{int}}[\phi]\right),$ (43)
@@
$S^{l}[\phi_{l}]=-\int_{k\leq\Lambda}\int dt\sqrt{-g}\left((\partial_{\mu}\phi_{l})^{2}+\mathcal{L}^{l}[\phi_{l}]\right),$ (46)
@@
$S=\int d^{3}x\int dt\sqrt{-g}\left(-\frac{1}{2}(\partial_{\mu}\phi_{l})^{2}-\frac{1}{2}(\partial_{\mu}\phi_{s})^{2}+\mathcal{L}[\phi_{l}]+\frac{\partial\mathcal{L}[\phi_{l}]}{\partial\phi_{l}}\phi^{s}-\frac{1}{2}\frac{\partial^{2}\mathcal{L}[\phi_{l}]}{\partial\phi_{l}^{2}}\phi_{s}^{2}\right).$ (47)
@@
$\Phi_{\mathrm{cl}}(k,t)=K(k,t)\phi_{\mathrm{cl}}+i\int dt^{\prime}G(k;t,t^{\prime})\frac{\delta S_{\mathrm{int}}}{\delta\Phi_{\mathrm{cl}}},$ (50)
@@
$(\Box-m^{2})G(k;t,t^{\prime})=\frac{i}{\sqrt{-g}}\delta(t-t^{\prime}),\qquad\lim_{t\to-\infty,t_{0}}G(k;t,t^{\prime})=0,$ (51)
@@
$G(k;t,t^{\prime})$ $=f^{*}_{k}(t)f_{k}(t^{\prime})\theta(t-t^{\prime})+f^{*}_{k}(t^{\prime})f_{k}(t)\theta(t^{\prime}-t)-\frac{f^{*}_{k}(t_{0})}{f_{k}(t_{0})}f_{k}(t)f_{k}(t^{\prime}).$ (52)
@@
$\rho_{\mathrm{red}}[\phi^{l}_{+},\phi^{l}_{-}]=\mathrm{Tr}_{k\geq\Lambda}\rho[\phi_{+},\phi_{-}].$ (53)
@@
$\Phi^{l}_{\mathrm{cl}}(k,t)=K^{l}(k,t)\phi^{l}_{\mathrm{cl}}+i\int dt^{\prime}G^{l}(k;t,t^{\prime})\frac{\delta S_{\mathrm{int}}}{\delta\Phi^{l}_{\mathrm{cl}}},$ (55)
@@
$-\frac{ig^{2}}{2}\int_{\bm{k}}\int d\eta\sqrt{-g}\int d\eta^{\prime}\sqrt{-g}\Phi_{-}^{l}(\eta)(G^{s})^{*}(k;\eta,\eta^{\prime})\Phi_{-}^{l}(\eta^{\prime}),$ (57)
@@
$\frac{K_{\sigma}(k,t)K^{*}_{\sigma}(k,t)}{2\mathrm{Re}\psi^{\sigma}_{2}}$ $=G_{+-}^{\sigma}(k;t,t^{\prime}).$ (61)
@@
$\left.+\frac{g^{2}}{2}\sum_{\pm}\int_{\bm{k}}\left(\int d\eta\sqrt{-g}\int d\eta^{\prime}\sqrt{-g}K_{\pm}^{l}(k,\eta)G^{s}_{\pm\pm}(k,\eta,\eta^{\prime})K^{l}_{\pm}(k,\eta)\right)\phi^{l}_{\pm}\phi^{l}_{\pm}+\mathcal{O}(g^{3})\right].$ (62)
@@
$\mathbf{\Sigma}=\left(\begin{array}[]{cc}\frac{1}{\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W}}&-\frac{\mathrm{Im}\psi_{l}+\mathrm{Im}\psi_{G}}{2(\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W})}\\ -\frac{\mathrm{Im}\psi_{l}+\mathrm{Im}\psi_{G}}{2(\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W})}&\frac{|\psi_{l}+\psi_{G}|^{2}-\psi_{W}^{2}}{4(\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W})}\end{array}\right).$ (68)
@@
$=\left(\frac{\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W}}{\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}+\psi_{W}}\right)^{1/2}.$ (69)
@@
$\gamma=1-\frac{\psi_{W}}{\mathrm{Re}\psi_{l}}+\mathcal{O}(g^{4}).$ (70)
@@
$=\frac{g^{2}}{2}\int d\eta\sqrt{-g}\int d\eta^{\prime}\sqrt{-g}G^{l}_{W}(k;\eta,\eta^{\prime})G^{s}_{W}(k;\eta,\eta^{\prime})+\mathcal{O}(g^{4}).$ (71)
@@ Appendix B Massless environment
$S=\int d^{4}x\sqrt{-g}\left[-\frac{1}{2}(\partial_{\mu}\phi)^{2}-\frac{1}{2}(\partial_{\mu}\sigma)^{2}-\frac{1}{2}m^{2}\sigma^{2}-\alpha\dot{\phi}\sigma\right].$ (72)
@@
$\langle\phi(\textbf{k})\phi(-\textbf{k})\rangle^{\prime}$ $=\frac{H^{2}}{2k^{3}}\left(1+\frac{\alpha^{2}}{H^{2}}\left((-2+\gamma_{E}+\log(-2k\eta_{0}))^{2}-\frac{\pi^{2}}{12}+1\right)+\mathcal{O}(k\eta_{0})\right),$ (75)
@@
$\langle\pi_{\phi}(\textbf{k})\pi_{\phi}(-\textbf{k})\rangle^{\prime}=\frac{k^{3}}{2H^{2}}\frac{1}{k^{2}\eta_{0}^{2}}\left(\frac{1}{2}-\frac{\alpha^{2}}{H^{2}}\left(-3+\gamma_{E}+\log(-2k\eta_{0})\right)^{2}+\mathcal{O}(-k\eta_{0})\right).$ (76)
@@
$\rho_{\phi}[\phi_{a},\phi_{s}]=\exp\left(-\frac{1}{2}\int_{\textbf{k}}\psi_{aa}\phi_{a}^{2}-\int_{\textbf{k}}\psi_{ar}\phi_{a}\phi_{s}-\frac{1}{2}\int_{\textbf{k}}\psi_{ss}\phi_{s}^{2}\right),$ (77)

### v2

@@ Acknowledgements
We acknowledge interesting conversations with Ahmed Almheiri, Cliff Burgess, Thomas Colas, Paolo Creminelli, Steve Gratton, Greg Kaplanek, Simon Lin, Juan Maldacena, Daniel Mata-Pacheco, Mehrdad Mirbabayi, Shoy Ouseph, Enrico Pajer, Pellegrino Piantadosi and Luis Zapata. We thank specially to Thomas Colas, Greg Kaplanek and Enrico Pajer for a careful reading of the manuscript. SC thanks the hospitality of New York University Abu Dhabi where part of this work was carried. SC is supported in part by the STFC Consolidated Grants ST/T000791/1 and ST/X000575/1 and by a Simons Investigator award 690508. The research of FQ is funded by a NYUAD research grant.
@@ Appendix A Details on the path integral derivations
$S[\phi]=\int d^{4}x\sqrt{-g}\left(-(\partial_{\mu}\phi)^{2}+\mathcal{L}_{\mathrm{int}}[\phi]\right)$ (43)
@@
$S^{l}[\phi_{l}]=-\int_{k\leq\Lambda}\int dt\sqrt{-g}\left((\partial_{\mu}\phi_{l})^{2}+\mathcal{L}^{l}[\phi_{l}]\right)$ (46)
@@
$S=\int d^{3}x\int dt\sqrt{-g}\left(-\frac{1}{2}(\partial_{\mu}\phi_{l})^{2}-\frac{1}{2}(\partial_{\mu}\phi_{s})^{2}+\mathcal{L}[\phi_{l}]+\frac{\partial\mathcal{L}[\phi_{l}]}{\partial\phi_{l}}\phi^{s}-\frac{1}{2}\frac{\partial^{2}\mathcal{L}[\phi_{l}]}{\partial\phi_{l}^{2}}\phi_{s}^{2}\right)$ (47)
@@
$\Phi_{\mathrm{cl}}(k,t)=K(k,t)\phi_{\mathrm{cl}}+i\int dt^{\prime}G(k;t,t^{\prime})\frac{\delta S_{\mathrm{int}}}{\delta\Phi_{\mathrm{cl}}}$ (50)
@@
$(\Box-m^{2})G(k;t,t^{\prime})=\frac{i}{\sqrt{-g}}\delta(t-t^{\prime}),\qquad\lim_{t\to-\infty,t_{0}}G(k;t,t^{\prime})=0$ (51)
@@
$G(k;t,t^{\prime})$ $=f^{*}_{k}(t)f_{k}(t^{\prime})\theta(t-t^{\prime})+f^{*}_{k}(t^{\prime})f_{k}(t)\theta(t^{\prime}-t)-\frac{f^{*}_{k}(t_{0})}{f_{k}(t_{0})}f_{k}(t)f_{k}(t^{\prime})$ (52)
@@
$\rho_{\mathrm{red}}[\phi^{l}_{+},\phi^{l}_{-}]=\mathrm{Tr}_{k\geq\Lambda}\rho[\phi_{+},\phi_{-}]$ (53)
@@
$\Phi^{l}_{\mathrm{cl}}(k,t)=K^{l}(k,t)\phi^{l}_{\mathrm{cl}}+i\int dt^{\prime}G^{l}(k;t,t^{\prime})\frac{\delta S_{\mathrm{int}}}{\delta\Phi^{l}_{\mathrm{cl}}}$ (55)
@@
$-\frac{ig^{2}}{2}\int_{\bm{k}}\int d\eta\sqrt{-g}\int d\eta^{\prime}\sqrt{-g}\Phi_{-}^{l}(\eta)(G^{s})^{*}(k;\eta,\eta^{\prime})\Phi_{-}^{l}(\eta^{\prime})$ (57)
@@
$\frac{K_{\sigma}(k,t)K^{*}_{\sigma}(k,t)}{2\mathrm{Re}\psi^{\sigma}_{2}}$ $=G_{+-}^{\sigma}(k;t,t^{\prime})$ (61)
@@
$\left.+\frac{g^{2}}{2}\sum_{\pm}\int_{\bm{k}}\left(\int d\eta\sqrt{-g}\int d\eta^{\prime}\sqrt{-g}K_{\pm}^{l}(k,\eta)G^{s}_{\pm\pm}(k,\eta,\eta^{\prime})K^{l}_{\pm}(k,\eta)\right)\phi^{l}_{\pm}\phi^{l}_{\pm}+\mathcal{O}(g^{3})\right]$ (62)
@@
$\Sigma=\left(\begin{array}[]{cc}\frac{1}{\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W}}&-\frac{\mathrm{Im}\psi_{l}+\mathrm{Im}\psi_{G}}{2(\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W})}\\ -\frac{\mathrm{Im}\psi_{l}+\mathrm{Im}\psi_{G}}{2(\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W})}&\frac{|\psi_{l}+\psi_{G}|^{2}-\psi_{W}^{2}}{4(\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W})}\end{array}\right),\qquad\mathrm{where}\ \psi^{(2)}=\psi_{l}+\psi_{G}$ (68)
@@
$=\left(\frac{\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}-\psi_{W}}{\mathrm{Re}\psi_{l}+\mathrm{Re}\psi_{G}+\psi_{W}}\right)^{1/2}$ (69)
@@
$\gamma=1-\frac{\psi_{W}}{\mathrm{Re}\psi_{l}}+\mathcal{O}(g^{4})$ (70)
@@
$=\frac{g^{2}}{2}\int d\eta\sqrt{-g}\int d\eta^{\prime}\sqrt{-g}G^{l}_{W}(k;\eta,\eta^{\prime})G^{s}_{W}(k;\eta,\eta^{\prime})+\mathcal{O}(g^{4})$ (71)
@@ Appendix B Massless environment
$S=\int d^{4}x\sqrt{-g}\left[-\frac{1}{2}(\partial_{\mu}\phi)^{2}-\frac{1}{2}(\partial_{\mu}\sigma)^{2}-\frac{1}{2}m^{2}\sigma^{2}-\alpha\dot{\phi}\sigma\right]$ (72)
@@
$\langle\phi(\textbf{k})\phi(-\textbf{k})\rangle^{\prime}$ $=\frac{H^{2}}{2k^{3}}\left(1+\frac{\alpha^{2}}{H^{2}}\left((-2+\gamma_{E}+\log(-2k\eta_{0}))^{2}-\frac{\pi^{2}}{12}+1\right)+\mathcal{O}(k\eta_{0})\right)$ (75)
@@
$\langle\pi_{\phi}(\textbf{k})\pi_{\phi}(-\textbf{k})\rangle^{\prime}=\frac{k^{3}}{2H^{2}}\frac{1}{k^{2}\eta_{0}^{2}}\left(\frac{1}{2}-\frac{\alpha^{2}}{H^{2}}\left(-3+\gamma_{E}+\log(-2k\eta_{0})\right)^{2}+\mathcal{O}(-k\eta_{0})\right)$ (76)
@@
$\rho_{\phi}[\phi_{a},\phi_{s}]=\exp\left(-\frac{1}{2}\int_{\textbf{k}}\psi_{aa}\phi_{a}^{2}-\int_{\textbf{k}}\psi_{ar}\phi_{a}\phi_{s}-\frac{1}{2}\int_{\textbf{k}}\psi_{ss}\phi_{s}^{2}\right)$ (77)